OpenStudy (anonymous):

John bought a new computer for $1,650. He paid a $160 down payment and financed the rest for 1 year at an interest rate of 7%. Find the total interest paid on the given amortized loan assuming that John makes monthly payments.

OpenStudy (anonymous):

choices are a)52.66 b)63.84 c)67.76 d)57.10

OpenStudy (tkhunny):

Basic Principles ALWAYS solves the problem - always. i = 0.07 -- Annual Interest Rate j = 0.07/12 = 0.0058333... -- Monthly Interest Rate r = 1+j = 1.0058333... -- Monthly Accumulation Factor v = 1/r = 0.99420 -- Monthly Discount Factor We have 12 payments, one month apart, and starting one month after purchase. 1650 - 160 = Pmt(v + v^2 + v^3 + ... + v^12) Do you know how to add up those 'v's?

OpenStudy (anonymous):

I don't know how to add the v's

OpenStudy (tkhunny):

You've probably seen it. Free demonstration of adding up finite geometric series: Call the sum something. 1) v + v^2 + v^3 + ... + v^12 = S 2) Multiply by the common ratio v^2 + v^3 + ... + v^12 + v^13 = Sv Subtract 2) from 1) v - v^13 = S - Sv Solve for S: \(S = \dfrac{v - v^{13}}{1-v}\) More simplifications will make it look more familiar, but this is it! 1650 - 160 = Pmt(v - v^13)/(1-v) Solve for Pmt.